OpenStudy (anonymous):

need help solving 0.4w=12 need to learn the steps

OpenStudy (anonymous):

i thought it was supose to be like 10(4w)=120?

OpenStudy (anonymous):

divide both sides by 0.4 0.4w 12 ----- = ----- 0.4 0.4 so w= 30

OpenStudy (anonymous):

ok thank you

OpenStudy (lgbasallote):

@rjmoore93 for the record what you're attempting is right too

OpenStudy (anonymous):

Divide by 0.4 both the sides..

OpenStudy (lgbasallote):

it helps to myltiply by 10 to get rid of decimals

OpenStudy (anonymous):

if the equation was longer and had more decimals will the rule apply to it too?

OpenStudy (anonymous):

If you do not know the decimal division, then you can get rid of the decimal by first multiplying both sides by 10(as you have done) and then divide both sides by 4

OpenStudy (anonymous):

give an example n ill explain..

OpenStudy (anonymous):

\[10 \times 0.4w = 10 \times 12 \implies 4w = 120\] Now divide by 4 both the sides..

OpenStudy (anonymous):

if there are multiple decimal numbers, then you see which one has the highest number of digits after the decimal point and multiply all accordingly

OpenStudy (anonymous):

for each number after the decimals depends on what i multiply it by like .12 would be 1,200?

OpenStudy (anonymous):

for removing decimal 0.2 is to be multiplied by 10 to get 2 0.57 is to be multiplied by 100 to get 57 2.478 is to be multiplied by 1000 to get 2478 and so on...

OpenStudy (anonymous):

so i move the decimal depending on how many zeros?

OpenStudy (anonymous):

look at it this way.... in 0.2 you have to shift the decimal towards right by 1 place to get 2 so multiply by one 10 in 0.57 you have to shift the decimal towards right by 2 places to get 57 so multiply by two 10s i.e. 10x10=100 in 2.478 you have to shift the decimal towards right by 3 places to get 2478 so multiply by three 10s i.e. 10x10x10 = 1000 Is it clear now???

OpenStudy (anonymous):

yea thank you I appreciate it.

OpenStudy (anonymous):

welcome...☺

OpenStudy (anonymous):

Change the coefficient of x to a fraction.\[\frac{4}{10}x=12\to \frac{2}{5}x=12\to 2x=60\to x=30 \]